OpenStudy (anonymous):

PLEASE HELP! REALLY LOST! The electric field in a 3.0mm×3.0mm square aluminum wire is 2.2×10^−2V/m. What is the current in the wire?

OpenStudy (anonymous):

Hi there. I think we can use current as the electric flow of charge per second. So let me present it: First, we take the distance as 3mm or 0.003m. Using the formula E=V/d, where d is distance, v voltage and E electric field strength, we make V the subject, being V=Ed or 2.2*10^-2*0.003=6.6*10^-5V Are you not given the force or charge?

OpenStudy (anonymous):

no, this is all i'm given. it's supposed to be I=. so i think i need an answer in amps. i tried just reversing J=I/A using the 2.2 as my J and the area, but that didn't work.

OpenStudy (irishboy123):

V = E•length, R = p • length/area, I = V/R = p • area / p. p (rho) for al is 2.65 x 10^-8.

OpenStudy (anonymous):

not really sure i follow that. up until I=V/R=....don't even know what R is supposed to be. e is electric field. but other than that....

OpenStudy (anonymous):

i'm ok up until I=V/R, then the equation seems all jumpy and weird. p/p and the p*area doesn't even match up with what you said V was supposed to be....

OpenStudy (irishboy123):

R is resistance. all you are given is a field, a x-sect area, and that it is Al wire. you need to calc volts from E • length, and resistance from rho•length/area. lengths cancel out.

OpenStudy (anonymous):

Current density is equal to the electric field times the conductivity. If the area of the wire is A then the current is J*A = conductivity * E*A look up the conductivity of Aluminum and you are home. note: conductivity is the inverse of resistivity